Night report 2021-04-02

Observers: Priscila Pessi, Ayan Mitra

Support: Phil Wiseman, Chris Frohmaier, Tomás Müller, Marcus Toy

TAT: Takashi Nagao, Lluís Galbany, Shubham Srivastav

EFOSC2 Calibrations:

Bias, dome_flats, arcs.

Skyflats but there were a lot of clouds so they might not be useful.

Weather:

Start: humidity 33% , wind 4.7 m/s, seeing 0.8

Middle: humidity 29%, wind 6.3 m/s, seeing 0.6

End: humidity 30%, wind 10.2 m/s, seeing 0.9

Comments:

The telescope closed at 7:30UT for one hour due to clouds.

We opened again at 8:28UT